# footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)

### v2

@@ 5.2. Motion of vertex positions
A proof of Theorem 5.9 is given in Appendix D.
@@
Since the curvature and torsion angles uniquely determine a curve up to a rigid transformation, Proposition 5.3 shows that the motion constructed in Theorem 5.1 is periodic up to a global rigid transform. Here, we control the global orientation by choosing the parameters $C$ and $\Gamma$ appropriately so that the motion becomes periodic in terms of vertex positions.
Proposition 5.11.
Let $l\in\mathbb{Z}$. If we set
(5.22) $C$ $=\frac{4\pi i}{y}\frac{1}{\Delta_{3}-\Delta_{1}},$
(5.23) $\Gamma$ $=-\frac{\pi}{y}\left(4r+2l+1\right),$
we have
(5.24) ${\gamma_{n}}(t+y)={\gamma_{n}}(t),$
for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$.
@@
First, we prove the periodicity for the $Z$-component. Denote the $Z$-component of $\gamma_{n}$ by ${\gamma_{n}}_{Z}$. From the pseudo-period of elliptic theta functions (A.12), we see
(5.25) $\begin{split}&{\gamma_{n}}_{Z}(t+y)-{\gamma_{n}}_{Z}(t)\\ &=-2\frac{\partial}{\partial z}\log\left\{\exp\left(\frac{C}{2}yz\right)\frac{\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it+iy\right)}{\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it\right)}\right\}\\ &=-2\frac{\partial}{\partial z}\log\exp\left\{\frac{C}{2}yz-\pi i\left(2iv\left(n-\frac{1}{2}\right)+\frac{2z}{\Delta_{3}-\Delta_{1}}+2it\right)+\pi y\right\}\\ &=-2\left(\frac{C}{2}y-\frac{2\pi i}{\Delta_{3}-\Delta_{1}}\right)=0.\end{split}$
For the $X$ and $Y$ components, from (3.10) we have to show $\frac{H_{n}(t+y)}{F_{n}(t+y)}\frac{F_{n}(t)}{H_{n}(t)}=1$.
If we write $P_{n}=iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it$, the left-hand side can be computed using (A.12) as
(5.26) $\begin{split}&\exp\left(i\Gamma y\right)\frac{\vartheta_{4}\left(P_{n}-2r+iy\right)\vartheta_{2}\left(P_{n}\right)}{\vartheta_{2}\left(P_{n}+iy\right)\vartheta_{4}\left(P_{n}-2r\right)}\\ &=-\exp\left(i\Gamma y\right)\frac{\exp\left\{-\pi i\left(2P_{n}-4r+iy\right)\right\}}{\exp\left\{-\pi i\left(2P_{n}+iy\right)\right\}}\\ &=-\exp\left\{i\left(\Gamma y+4\pi r\right)\right\}=1.\end{split}$
5.3. Evolution of curvature
The curve constructed in Theorem 5.1 reveals an interesting recurrence equation for $\kappa$.
@@ 5.3. Evolution of curvature
A proof of Proposition 5.12 is given in Appendix E. The evolution of the signed curvature angle satisfies the semi-discrete mKdV equation in Proposition 5.7.
@@
(5.28) $\dot{\kappa}_{n}=2\beta^{2}\eta_{3}\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right),$
where
(5.29) $\eta_{3}=i\frac{\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}{\vartheta_{1}\left(2iv\right)\vartheta_{3}\left(0\right)}\left(d_{1}(iv)-d_{3}(iv)\right).$
Combined with Proposition 5.7, (5.6), and the fact that $e^{i\lambda}\delta_{2}\in\mathbb{R}$, we see
(5.30) $2\beta^{2}\eta_{3}=\left(2\beta\beta_{1}\cos\lambda+2\beta\beta_{2}\sin\lambda\right).$
A proof of Theorem 5.13 is given in Appendix F.
@@ 5.4. Potential function of curvature
For fixed $v,y$, define the function $\Theta$ as follows.

### v3

@@ 5.2. Motion of vertex positions
Proof.
We determine the necessary and sufficient condition for $C$ and $\Gamma$ to satisfy
(5.112) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,$
(5.113) $\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\rho^{2},$
where $\rho\in\mathbb{R}$ is a constant. From (5.83) and (5.104) we see
(5.114) $\begin{split}2iF_{n}^{3}\widetilde{G}_{n}\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=&\left(P_{n}-P_{n}^{*}\right)\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\right)\\ &+\left(P_{n}+P_{n}^{*}\right)\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ &+2\beta_{2}|G_{n}|^{2}\left\{\left(\beta_{6}+C\right)F_{n}^{2}+\beta_{7}|G_{n}|^{2}\right\}\\ =&2F_{n}^{2}|G_{n}|^{2}\left\{2\eta_{1}\left(2\beta_{3}-\beta_{1}\right)\beta_{2}+2i\eta_{2}+\beta_{2}\left(\beta_{6}+C\right)\right\}\\ &+2\beta_{2}|G_{n}|^{4}\left\{2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}+\beta_{7}\right\}\\ =&2F_{n}^{2}|G_{n}|^{2}\left\{2\eta_{1}\left(2\beta_{3}-\beta_{1}\right)\beta_{2}+2i\eta_{2}+\beta_{2}\left(\beta_{6}+C\right)\right\},\end{split}$
where we have used (5.85). Therefore, we obtain the following as a necessary and sufficient condition for (5.112) to hold.
(5.115) $2\eta_{1}\left(2\beta_{3}-\beta_{1}\right)\beta_{2}+2i\eta_{2}+\beta_{2}\left(\beta_{6}+C\right)=0,\\$
Using (5.84), we rewrite this in terms of $C$ and $\Gamma$ to obtain
(5.116) $i\Gamma\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)+\beta_{6}+C+\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0.$
Next is the calculation of (5.113). The following holds
(5.117) $\begin{split}&F_{n}^{4}\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle\\ =&\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\right)^{2}\\ &-\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)^{2}\\ &+\left\{\left(\beta_{6}+C\right)F_{n}^{2}+\beta_{7}|G_{n}|^{2}\right\}^{2}\\ =&F_{n}^{4}\left\{16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}\right\}\\ &+F_{n}^{2}|G_{n}|^{2}\left\{4\eta_{1}^{2}\left(1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right)-8i\eta_{1}\eta_{2}\beta_{2}\left(2\beta_{3}-\beta_{1}\right)+4\eta_{2}^{2}+2\beta_{7}\left(\beta_{6}+C\right)\right\}\\ &+|G_{n}|^{4}\left(-4\eta_{1}^{2}\left(\beta_{4}+\beta_{4}^{*}\right)^{2}-8i\eta_{1}\eta_{2}\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)+4\eta_{2}^{2}\beta_{2}^{2}+\beta_{7}^{2}\right)\\ =&F_{n}^{4}\left\{16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}\right\}\\ &+F_{n}^{2}|G_{n}|^{2}\left\{4\eta_{1}^{2}\left(1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right)-8i\eta_{1}\eta_{2}\beta_{2}\left(2\beta_{3}-\beta_{1}\right)+4\eta_{2}^{2}+2\beta_{7}\left(\beta_{6}+C\right)\right\},\end{split}$
where we have used (5.86). Therefore, we obtain the following as a necessary and sufficient condition for (5.113) to hold.
(5.118) $4\eta_{1}^{2}\left(1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right)-8i\eta_{1}\eta_{2}\beta_{2}\left(2\beta_{3}-\beta_{1}\right)+4\eta_{2}^{2}+2\beta_{7}\left(\beta_{6}+C\right)=0.$
By using (5.85) and (5.116), (5.118) can be written in terms of $\Gamma$ as follows.
(5.119) $i\beta_{4}\beta_{4}^{*}\left(\beta_{4}-\beta_{4}^{*}\right)\Gamma^{2}+2\beta_{4}\beta_{4}^{*}\left(\beta_{5}+\beta_{5}^{*}\right)\Gamma+i\left(\beta_{4}{\beta_{5}^{*}}^{2}-\beta_{4}^{*}\beta_{5}^{2}\right)=0.$
This is a quadratic equation in $\Gamma$ and has the following solutions.
(5.120) $\Gamma_{\pm}=\frac{-\beta_{4}\beta_{4}^{*}\left(\beta_{5}+\beta_{5}^{*}\right)\pm\sqrt{\beta_{4}\beta_{4}^{*}\left(\beta_{4}\beta_{5}^{*}+\beta_{4}^{*}\beta_{5}\right)^{2}}}{i\beta_{4}\beta_{4}^{*}\left(\beta_{4}-\beta_{4}^{*}\right)}.$
Since
(5.121) $\begin{split}&\beta_{4}\beta_{4}^{*}\left(\beta_{4}\beta_{5}^{*}+\beta_{4}^{*}\beta_{5}\right)^{2}=\left(\beta_{4}\beta_{4}^{*}\right)^{2}\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2},\\ &\beta_{4}-\beta_{4}^{*}=\frac{\vartheta_{1}(2r)\vartheta_{3}(0)}{\vartheta_{3}(2r)\vartheta_{1}(iv)},\\ &\beta_{5}+\beta_{5}^{*}=-i\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)}{\vartheta_{3}(2r)\vartheta_{1}(iv)},\end{split}$
holds, $\Gamma_{\pm}$ can be rewritten as
(5.122) $\Gamma_{\pm}=\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)\pm\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)}.$
Substituting this into (5.116) yields the expression for $C_{\pm}$ as well. From the above calculations, we see that when $C_{\pm}$ and $\Gamma_{\pm}$ are chosen as in (5.109), the following holds
(5.123) $\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}.$
We calculate this to compute $\rho^{2}$. From (5.84), we see that when $(C,\Gamma)=(C_{+},\Gamma_{+})$, the following holds
(5.124) $16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)=4\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}.$
From (5.115) and (5.85), we see that when $(C,\Gamma)=(C_{+},\Gamma_{+})$, the following holds
(5.125) $\left(\beta_{6}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}\frac{\left(\vartheta_{3}(r)^{4}-\vartheta_{1}(r)^{4}\right)^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}.$
Adding the above equations yields
(5.126) $\rho^{2}=16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}=\rho_{+}^{2},$
where we have used
(5.127) $\vartheta_{3}(r)^{4}+\vartheta_{1}(r)^{4}=\vartheta_{3}(2r)\vartheta_{3}(0)^{3}.$
In the same way, we see that when $(C,\Gamma)=(C_{-},\Gamma_{-})$,
(5.128) $\rho^{2}=16\eta_{1}^{2}\left(\beta_{1}-\beta_{3}\right)\beta_{3}+\left(\beta_{6}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{4}(r)^{4}\vartheta_{2}(r)^{4}}=\rho_{-}^{2},$
holds. ∎
@@
5.3. Evolution of curvature
The curve constructed in Theorem 5.1 reveals an interesting recurrence equation for $\kappa$. To obtain the recurrence equation, we use the following Lemma.
Lemma 5.13.
The $\tau$ functions (5.1) satisfy the following equations.
(5.129) $\widetilde{G}_{n}^{2}+\beta_{1}^{2}F_{n}^{2}=F_{n+1}F_{n-1},$
(5.130) $\beta_{1}^{2}\widetilde{G}_{n}^{2}-\frac{\vartheta_{1}\left(iv\right)^{4}}{\vartheta_{3}\left(iv\right)^{4}}\beta_{1}^{4}F_{n}^{2}=\widetilde{G}_{n+1}\widetilde{G}_{n-1}.$
@@ 5.3. Evolution of curvature
Using (3.1)-(3.3), we see that (5.129) hold. Putting
[…]
in (A.10), we obtain (5.130). ∎
@@
Proof.
Using (5.129) and (5.130), we see that (5.136) holds. ∎
The evolution of the signed curvature angle satisfies the semi-discrete mKdV equation in Proposition 5.9. To obtain the semi-discrete mKdV equation, we use the following Lemma.
Lemma 5.15.
The $\tau$ functions (5.1) satisfy the following equation.
(5.137) $D_{t}\widetilde{G}_{n}\cdot F_{n}=\beta_{8}\left(\widetilde{G}_{n+1}F_{n-1}-\widetilde{G}_{n-1}F_{n+1}\right),$
where
(5.138) $\beta_{8}=i\frac{\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}{\vartheta_{1}\left(2iv\right)\vartheta_{3}\left(0\right)}\left(d_{1}(iv;y)-d_{3}(iv;y)\right).$
Proof.
After differentiating the equation (A.3) with
[…]
by $h$ and putting $h=0$, we obtain
(5.144) $\begin{split}&D_{t}\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\cdot\left\{\vartheta_{3}(iv)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\\ &=\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(\mu_{n-1}-\frac{1}{2}iv+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\right\}\\ &-\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(\mu_{n-1}-\frac{1}{2}iv+it\right)\right\}.\end{split}$
After rearranging (5.144), we obtain (5.137). ∎
@@
(5.145) $\dot{\kappa}_{n}=2\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
Proof.
Differentiating the signed curvature angle yields the following.
(5.146) $\dot{\kappa}_{n}=2\frac{d}{dt}\arctan\left(\frac{\widetilde{G}_{n}}{\beta_{1}F_{n}}\right)=2\beta_{1}\frac{D_{t}\widetilde{G}_{n}\cdot F_{n}}{\widetilde{G}_{n}^{2}+\beta_{1}^{2}F_{n}^{2}}.$
Using (5.129) and (5.137), we see that the signed curvature angle satisfies
(5.147) $\dot{\kappa}_{n}=2\beta_{1}^{2}\beta_{8}\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
Since
(5.148) $\beta_{1}\beta_{8}=\eta_{1}\cos\lambda+\eta_{2}\sin\lambda,$
holds, this is nothing but the semi-discrete mKdV equation (5.145). ∎
@@ 5.4. Potential function of curvature
For fixed $v$ and $y$, define the function $\Theta$ as follows.